OpenStudy (anonymous):

If a runner burns 30 calories in 5 hours and 120 calories in 20 hours, how many calories is he burning per minute?

OpenStudy (anonymous):

How many minutes are there in 5 hours?

OpenStudy (anonymous):

300

OpenStudy (a_clan):

30 cal in 5 hrs = 6 cal in 1 hr(60 min) = 6/60 cal in 1 min

OpenStudy (anonymous):

So that's 30 calories in 300 minutes. Or 30/300=0.1 calories per minute

OpenStudy (anonymous):

I could swear I put in both 1 calorie and 0.1 calorie and it said it was wrong.

OpenStudy (katrinakaif):

Well first we need to figure how many mins we have in 5 hours. We know that in: 1hr = 60mins. So to figure out how many we have in 5 mins, we use 60(5) = 300 as you said. Next set up a proportion: 300/30 = 60/x 300 = number of mins in 5 hours burns '30' calories and 60 mins = 1 hr burns x calories Cross Mulitply and get: 300x = 1800 Divide by 300 onhn both sides giving..? 6 calories per min :)

jhonyy9 (jhonyy9):

6*5=30 6*20=120 - from this resulted that he burning 6 calories per minute

OpenStudy (rosey):

runners r hot <3